2)   If the break-even point is 15,000 units, the selling price is $95, and the unit variable cost is $75, what are the company's total fixed costs? (3 mark)
3)   Furry Friends food manufacturing has compiled the following estimates for operations.
Sales: $762,000
Fixed cost: $150,000
Total variable cost: $590,000   
a)   Find the total cost (3 mark)
b)   Find the net income. (3 mark)
Answer #1

2. Breakeven qty = 15000

Selling price = 95

Total revenue = 95 * 15000 = 1425000

Total cost = Variable cost + Fixed cost = 75 * 15000 + Fixed cost

As per condition given

75 * 15000 + Fixed cost = 1425000

Fixed cost = 1425000 - 75 * 15000 = 300000

3.

a. Total cost = FC + VC = 150000 + 590000 = 740000

b. Net income = Sales - Total cost = 762000 - 740000 = 22000